OpenStudy (anonymous):

Differentiate  y=(1+2x^2)cot^-1 2x

OpenStudy (tkhunny):

Clarify \(y = (1+2x^2)\cdot \cot^{-1}(2x)\) ?? Product Rule Chain Rule Go!

OpenStudy (anonymous):

YES !

OpenStudy (tkhunny):

Chop it up. Write it out. \(\dfrac{dy}{dx} = (1+2x^{2})\cdot \left[\dfrac{d}{dx}\cot^{-1}(2x)\right] + \cot^{-1}(2x)\cdot \left[\dfrac{d}{dx}(1+2x^{2})\right]\) Now what?

OpenStudy (anonymous):

Thanks !!

OpenStudy (anonymous):

I also need help with this question as well Find the derivative \[y=arcoss \frac{ b+acosx }{ a+bcosx} , 0\le x \le \pi , a > b >0 \]

OpenStudy (tkhunny):

Chain Rule Quotient Rule Go! Let's see your best work.